OpenStudy (anonymous):

What are the roots of -12x^2 + 11x − 3 = 0?

OpenStudy (anonymous):

Hi To find the roots/solutions/zeros of this quadratic equation you have to either use the quadratic formula or factor it. quadratic formula is : \(\LARGE x\; =\; \frac {-b\; \pm\; \sqrt {b^2\; -\; 4ac}}{2a} \)

OpenStudy (anonymous):

wats would the be pluged in idk where to put the numbers

OpenStudy (anonymous):

\(ax^2+bx+c=0\) so in your case, a=-12, b=11 and c=-3